OpenStudy (anonymous):

Two factory plants are making TV panels. Yesterday, Plant A produced 7000 fewer panels than Plant B did. Four percent of the panels from Plant A and 2% of the panels from Plant B were defective. How many panels did Plant B produce, if the two plants together produced 680 defective panels?

OpenStudy (anonymous):

suppose botany is about "plants". Production from plant a = (b-4000) 3% of a plus 5% of b = 1320 so, (b-4000) x 3/100 + b x 5/100 = 1320 From this, you can work out b. Then you can get a.